• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:03:36 on November 8, 1998

    Entry number 4920

    keyword=Beam Time Accounting
    Hall A

    11/8/1998 0800-1600
    Experiment #: E94-010
    Shift Leader: B Hersman
    Shift Workers: S Incerti, S Malov, R Sulieman

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    1 1 1 1 1 1 1 1 8
    Beam Available
    but Not in Use
    0 0 0 0 0 0 0 0 0
    Beam Not Available
    or Unacceptable
    0 0 0 0 0 0 0 0 0
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 1 1 1 1 1 1 8
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: Beam stable at 10 uA, increased at ~2:00 to 15 uA Target polarization stable (slightly rising) at 32% (two measurements)netscape Final NMR measurement at 4:00 will determine if 15 uA causes significant depolarization during final 2 hours.

    Data Entry by: Bill Hersman